OpenStudy (anonymous):

School books are printed on a four color web offset printing press a cylinder on a press has a 13.37 inch diameter. The linear speed of a point on the cylinder's surface is 16.53 feet per second. What is the angular speed of the cylinder in revolutions per hour? Printers often refer to the angular speed as impressions per hour or IPH.

OpenStudy (anonymous):

School books are printed on a four color web offset printing press a cylinder on a press has a 13.37 inch diameter. The linear speed of a point on the cylinder's surface is 16.53 feet per second. What is the angular speed of the cylinder in revolutions per hour? Printers often refer to the angular speed as impressions per hour or IPH. HELP!

OpenStudy (perl):

we can use the equation linear speed = radius * angular speed, where angular speed is in radians/ second

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

C=13.37 pi we know that

OpenStudy (perl):

16.53 = (13.37 /2) * w

OpenStudy (perl):

w = 16.53 / ( 13.37 / 2 ) rad/second

OpenStudy (perl):

now we want to change that to revolution per hour

OpenStudy (anonymous):

yea but we're supposed to set up ratios in our work... so would it be 16.53 ft/ 1sec x 2/37 ?

OpenStudy (anonymous):

sorry.. kinda confused can u like draw it

OpenStudy (perl):

drawing doesnt work

OpenStudy (perl):

w = (2 * 16.53 ) / 13.37 * rad/sec * rev/ (2pi * rad) * 3600 sec/ 1hr

OpenStudy (anonymous):

why not? sorry im just trying to visualize this

OpenStudy (perl):

not working on my computer?

OpenStudy (anonymous):

... ugh mine neither

OpenStudy (anonymous):

ok so thats it?

OpenStudy (anonymous):

but they say its 16.53 ft/sec so why multiply by 2? and whats the units? is it just 33.06 ft / 13.37 inches?

OpenStudy (perl):

because 1/(1/2) = 2

OpenStudy (perl):

do you agree that v = r* w , where v is linear velocity , r is radius, and w is angular velocity and units of w are rad/sec

OpenStudy (anonymous):

yup

OpenStudy (perl):

ok